OpenStudy (anonymous):

Write the sum or difference in the standard form a + bi. ( 3 + 3i) - ( -6 + i)

OpenStudy (owlcoffee):

When working with complex numbers, we shall make the difference: (3+3i)-(-6+i) = (3-6)-(3+1)i since complex numbers are treated as vectors, and we remember that in the substraction of two vectors, we need do the sum with the opposite: \[(3-6)+(-3-1)i\] so the result is: \[-3-4i\]

OpenStudy (anonymous):

I see what you did but the choices for the question were -9-2i, -3+4i, 9-2i, 9+2i

OpenStudy (owlcoffee):

Oh, do the opposite of the second before grouping them, try that.

OpenStudy (anonymous):

(3+3i)-(-6+i) 3+6+3i-i=9+2i you shoud add imaginary form with each other and real form with eachother too

OpenStudy (anonymous):

thank you both!

OpenStudy (anonymous):

your welcome